OpenStudy (lgbasallote):

\[\int \frac{1+x}{1+x^2}dx\] how??

OpenStudy (lgbasallote):

uhmm i think i agree with that...

OpenStudy (anonymous):

use partial fraction decmpostion

OpenStudy (lgbasallote):

hmm \[\frac{1+x}{1+x^2} = \frac{Ax + B}{1+ x^2}?\] i dont think thatll work :/

OpenStudy (anonymous):

1/(1+x^2) + x/(1+x^2) easy integration

OpenStudy (lgbasallote):

what did you do? o.O

OpenStudy (lgbasallote):

ohhh

OpenStudy (lgbasallote):

but 1/1+x^2 doesnt seem easy integration :p

OpenStudy (lalaly):

Callisto is awesome

OpenStudy (callisto):

use trigo sub...

OpenStudy (lalaly):

1/1+x^2 is always tan^-1(x)

OpenStudy (lgbasallote):

can i do this by u-sub?

OpenStudy (lgbasallote):

ALWASY?

OpenStudy (anonymous):

its arctgx

OpenStudy (lalaly):

the integral of 1/1+x^2 i mean is always tan^-1(x) and in the second integral u can use u subst.

OpenStudy (anonymous):

i mean: Integral of 1/1+x^2 = arctg(x) +c

OpenStudy (lgbasallote):

but can i use u sub from the start

OpenStudy (lalaly):

yes u can do that

OpenStudy (lgbasallote):

how??

OpenStudy (lalaly):

lol im not sure but ill try

OpenStudy (lgbasallote):

this is supposed to be answered by u sub i think

OpenStudy (anonymous):

i don't think so

OpenStudy (lgbasallote):

no?

OpenStudy (lalaly):

let u=1+x^2 du=2x

OpenStudy (anonymous):

arctg(x) +1/2Ln(1+x^2) +c

OpenStudy (lgbasallote):

i dont need the answer -_- im looking for the u sub method

OpenStudy (lalaly):

du=2xdx \[\int\limits{(\frac{1}{1+x^2}+\frac{x}{1+x^2}})dx\]first one is tan^-2(x) second one \[\frac{1}{2} \int\limits{\frac{du}{u}}\]

OpenStudy (lgbasallote):

lol you still used trig sub :/

OpenStudy (lgbasallote):

or whatever that's called

OpenStudy (lgbasallote):

is it impossible? :/

OpenStudy (anonymous):

there is a general wayto integrate the expression: \[Mx+N/x ^{2}+px+q\] where it's supoused that denominator has no real roots

OpenStudy (lgbasallote):

ok...that just became complicated o.O

OpenStudy (anonymous):

but that's what you whant, no?

OpenStudy (lgbasallote):

nevermind..ill just go with that solution -_-

OpenStudy (anonymous):

if you whant i explain you the method, but it will take for a long time i guess

OpenStudy (anonymous):

break this down to: \[\int\limits (1/x^2+1) + (x/x^2+1) dx\] \[=\tan^{-1}x+\int\limits x/x^2+1dx\] let u=x^2+1 du=2xdx du/2=dx \[=\tan^{-1}x+1/2\int\limits1/udu=\tan^{-1}x+\ln\left| u \right|/2+c\] \[=\tan^{-1}x+(\ln\left| x^2+1 \right|)/2+c\]

OpenStudy (australopithecus):

\[\int\limits_{}^{} (1+x)dx/(1+x^{2}) = \int\limits_{}^{} dx/(1+x^{2}) + \int\limits_{}^{} xdx/(1+x^{2})\] for the one integral use u sub u = 1 + x^(2) du = 2xdx du/2 = xd thus \[(1/2) \int\limits_{}^{} du/(u) + \int\limits_{}^{} dx/(1+x^{2}) = (\ln|1+x^{2}|/2) + \int\limits_{}^{} dx/(1+x^{2})\] for the second integral we use trigonometric substitution tan^(2)(x) + 1 = sec^(2)(x) thus x = tan(theta) theta = arctan(x) now we need to write it in terms of theta so d(theta) sec^(2)(theta) = dx do substitution \[\int\limits_{}^{} d(\theta)\sec^{2}(\theta)/1 + \tan^{2}(\theta) = \int\limits_{}^{} d(\theta)\sec^{2}/\sec^{2}(\theta) = \int\limits_{}^{} d(\theta) = \theta\] since theta = arctan(x) we sub that into the answer and we get (ln|1+x^(2)|/2) + arctan(x) + c